a blue print has a scale of 1 in, representing 45ft. How long is a side of a buliding represented by 2 1/2 in. ?
If it takes 10 minutes to saw a long into 3 pieces, how long does it take to saw it into 4 pieces?
The ratio of boys to girls in a class is 8:7. If the class has 60 students, how many are girls?

Answer by carl(7)

The best way is to ratio these problems, let's get started.....
1 inches : 45 feet = 2.5 (2 1/2) inches : X feet
2.5 times 45 = 112.5 / 1 = 112.5 feet
-------------------------------------------------
10 mins : 3 pcs = X mins : 4 pcs
4 times 10 = 40 / 3 = 13.3 mins.
------------------------------------------------
7 boys : 8 girls...... total 60 students....
*** Think 30 boys and 30 girls would be a 1 to 1 ratio... but you need a 7 to 8 ratio ****
SO.................
7 X 4 = 28 boy students
8 X 4 = 32 girl students
28 + 32 = 60 total students
So..... the answer is 32 girls.
------------------------------------------------
Good luck!
Carl